# Back on Track: Bundle Adjustment for Dynamic Scene Reconstruction

### v1

@@ 3.3 Stage III: Global Refinement
Bundle adjustment recovers the camera poses and adjusts the depths of the query points in their source frames of reference. However, BA processes only a sparse set of query points, thus the 3D positions of the other points in the depth maps remain unaffected. To ensure consistency of the depth maps with the sparse set of accurate 3D tracks estimated by BA, we need global refinement. Towards this goal, we introduce a function $\mathcal{H}_{\theta}:D\rightarrow\hat{D}$, where $\hat{D}$ is the refined depth map and $\theta$ is the parameter set. Global refinement jointly optimizes the function parameters $\theta$ and sparse 3D trajectories to achieve consistent geometry.
@@
Inspired by previous work [21], we define a 2D scale grid of resolution $H_{g}\times W_{g}$ for each frame, $\theta_{t}\in\mathbb{R}^{H_{g}\times W_{g}}$. The grid has a coarser resolution compared to the original image resolution $H\times W$. Its goal is to scale the depth estimates $D_{t}$: For each 2D query point $\mathbf{x}$ in frame $t$, we determine the associated depth scale using bilinear sampling on the grid $\theta_{t}$. The refined depth is then
@@
is the 3D position derived from back-projecting 2D points using the refined dense depth. The weight $W_{\text{static}}$ filters out the dynamic points as $W_{\text{static}}=(1-m^{i}_{a})\cdot(1-m^{i}_{b}).$ Intuitively, $\mathcal{L}_{\text{rigid}}$ facilitates constant distances between arbitrary static points $a$ and $b$ in the target frame $j$ and the source frame $i$. We optimize the variables using the Adam optimizer [20].
Figure 5: Qualitative camera pose estimation results on Sintel [6], Shibuya [30] and Epic Fields [40]. Visualizations demonstrate that our method achieves more robust and accurate camera trajectories in challenging dynamic scenes.
@@ 4.1 Implementation Details
We train our 3D tracker on the TAP-Vid-Kubric training set [10] consisting of 11,000 sequences, each with 24 frames. The training process uses AdamW optimizer [24] with a learning rate of $3\times 10^{-4}$ for a total of 100,000 steps performed on 4 NVIDIA A100 GPUs. During training, we set the weights for visibility and dynamic labels to $w_{1}=5$ and $w_{2}=5$, employ $K=4$ iterative updates, a window size of $S=12$, and a total of $N=512$ query points. For BA, we use a window of 15 frames and apply 4 Gauss-Newton updates per window. We set $\alpha=0.05$ to balance reprojection and depth loss in  Eq. 8. For BA weight filtering we set $\delta_{v}=0.9$ and $\delta_{m}=0.9$. For further implementation details, please refer to Appendix A.
@@ 4.2 Camera Pose Evaluation
Datasets.   Three datasets are used for camera pose evaluation. MPI Sintel [6] provides sequences from 3D animated films featuring complex, fast object motion and exposure effects like motion blur. We use MPI Sintel to evaluate dynamic VO performance, with sequences spanning 20 to 50 frames. AirDOS Shibuya [30] includes sequences of 100 frames tracking over 30 individuals in two scene types. “Road Crossing” features diverse object motion, while “Standing Humans” contains dominant camera motion. EPIC Fields [40] includes egocentric videos with dynamic interactions like moving hands with kitchen tools, therefore is ideal for studying long dynamic scenarios. We randomly select 8 sequences with reliable ground-truth trajectories, sample every third of the first 3000 frames, yielding 1000 frames per sequence.
Method MPI Sintel [6] AirDOS Shibuya [30] Bonn [28]
@@
Robust-CVD [21] 0.703 47.8 - - - -
CasualSAM [47] 0.387 54.7 0.261 79.9 0.169 73.7
MonST3R [18] 0.335 58.5 (0.208) (71.2) 0.063 96.4
ZoeDepth [3] 0.467 47.3 0.571 43.8 0.087 94.8
@@
test-time network finetuning; depth from image pairs; the numbers in ()($\cdot$)( ⋅ ) denote results on a sub-sequence due to memory constraints.
Table 2: Depth evaluation results on Sintel [6], Shibuya [30], and Bonn [28] datasets. Our method achieves improved depth accuracy compared to the depth priors from ZoeDepth, demonstrating the effectiveness of our depth refinement.
Metrics.   Three metrics are adopted. Absolute Translation Error (ATE) calculates the root mean square error between the estimated and the ground-truth trajectories. Relative Translation Error (RTE) and Relative Rotation Error (RRE) measure translation (in meters) and rotation (in degrees) errors over a set distance, with both metrics averaged across all poses. We perform all evaluations after Sim(3) Umeyama alignment [41] with the ground-truth trajectory.
Results.   We compare BA-Track against learning-based SLAM and VO approaches, such DROID-SLAM [37], ParticleSfM [48], DytanVO [34], and LEAP-VO [7]. Additionally, we compare BA-Track to test-time optimization techniques, such as Robust-CVD [21] and CasualSAM [47], which typically require several hours of processing time.
Tab. 1 shows a quantitative comparison of our method against state-of-the-art approaches across three benchmarks. Our approach consistently delivers superior performance in ATE and remains competitive in RTE and RRE. On the MPI Sintel dataset, our method significantly outperforms all baselines. As Fig. 5 illustrates, BA-Track produces notably more accurate trajectory estimates, particularly in dynamic scenes where other methods struggle. On the challenging Epic Fields dataset, characterized by rapid dynamic content and complex camera motion, BA-Track maintains compelling VO accuracy. Overall, BA-Track demonstrates remarkable accuracy across all benchmarks with high memory efficiency. By contrast, prior work like MonST3R can only process up to 90 frames at once on a 48GB GPU.
@@ 4.3 Depth Evaluation
Datasets.   MPI Sintel [6] and AirDOS Shibuya [30] are selected as both provide dense ground-truth depth maps. We add a real-world dataset, Bonn RGB-D, which captures indoor activities such as object manipulation across 24 dynamic sequences. Following Zhang et al. [46], we select 5 dynamic sequences, each consisting of 110 frames.
Metrics.   We adopt the absolute relative error (Abs Rel) and the threshold accuracy (TA) $\delta<1.25$. Let $d_{i}$ and $d_{i}^{*}$ denote the predicted and the ground-truth depths for pixel $i$. Abs Rel is $\frac{1}{N}\sum_{i=1}^{N}\frac{|d_{i}^{*}-d_{i}|}{d_{i}^{*}}$ with $N$ the total number of pixels. TA is computed as the percentage of pixels satisfying $\max\left(\frac{d_{i}^{*}}{d_{i}},\frac{d_{i}}{d_{i}^{*}}\right)<1.25$. We apply shift-scale alignment by computing a single shift and scale factor over the entire video to align the predicted depths with the ground truth [46].
Figure 6: Motion decoupling on the DAVIS dataset [19]. From left to right, the three columns show the reference frame for queries, total observed motion, and the static component. Green indicates static trajectories, and red indicates dynamic ones. The static component retains points on dynamic objects in their original (static) positions within the corresponding reference frame.
Results.   We compare BA-Track to joint pose-depth optimization methods in Tab. 2. Using ZoeDepth [3] as input, our method consistently enhances depth estimates and achieves competitive accuracy compared to other joint refinement methods. As we show later in Sec. 4.5, our global refinement plays a crucial role here, allowing BA-Track to handle challenging scenarios with dynamic content. On MPI Sintel, we achieve an Abs Rel of 0.408 and an accuracy of 54.1%, demonstrating comparable depth quality w.r.t. leading methods trained on much larger datasets, such as MonST3R [46]. In more complex scenarios, such as the outdoor urban scenes in AirDOS Shibuya, our method maintains robust results with an Abs Rel of 0.299 and 55.1% accuracy, outperforming competitors, such as RobustCVD [21]. Our scale-map-based refinement offers efficient optimization with few parameters while remaining relatively simple. Investigating more advanced deformation models, such as neural networks, could be an interesting direction for future work.
@@ 4.4 Motion Decoupling
In our 3D tracker, motion decoupling isolates the static (camera-induced) component from the observed point motion, allowing us to treat dynamic points as if they were stationary. This approach reduces the impact of dynamic regions on camera pose estimation and dramatically enhances the robustness of BA. It also aids the subsequent depth refinement, providing more coherent reconstructions of both static and dynamic parts. Fig. 6 illustrates the observed and static point motion on the DAVIS [19] dataset. Red points indicate dynamic trajectories, while green points represent static ones, highlighting the 3D tracker’s ability to decouple the dynamic component from the observed motion.
To further demonstrate the benefits of motion decoupling, we analyze alternative motion representations for BA optimization in Tab. 3. By applying motion decoupling with the dynamic tracker $\mathcal{T}_{\text{dyn}}$, we achieve a dramatic performance boost: The ATE drops by nearly a half, from 0.137 (a) to 0.065 (e). We observe a further moderate boost in the VO accuracy by setting the weights of dynamic points in BA optimization to zero (f). Note that it surpasses the VO accuracy of the setting with a single tracker predicting the observed motion (b). As an alternative tracker architecture, we investigate an approach with a static tracker (c-d). In this case, we train the tracker to directly regress the static component. However, this approach leads to inferior VO accuracy compared to our decoupling approach (e-f). Presumably, a single network struggles to learn two aspects simultaneously: visual tracking and motion patterns. In contrast, our dual-network design addresses these tasks independently.
@@ 4.5 Dynamic Reconstruction
Table 3: Ablation study of dynamic handling methods on Sintel [6]. Combining motion decoupling with camera masking achieves the best results. * denotes a setting trained for ablation.
Method Depth Refinement Bonn crowd2 [28]
@@
Table 4: Ablation study on depth refinement on Bonn crowd2 sequence [28]. Applying both losses has a complementary effect.
We present qualitative results for camera tracking and 3D reconstruction on dynamic scenes in Fig. 2 using an alternative depth model from UniDepth [29], illustrating that our model is agnostic to the choice of depth priors. Using estimated camera poses, we back-project points into 3D via refined depth maps, yielding accurate camera motion estimation and consistent reconstructions. Our method remains robust in challenging scenarios with multiple moving people, rapid egocentric camera motion, and dynamic object interactions.
The superior reconstruction quality primarily stems from the proposed global refinement. To demonstrate this, we conduct an ablation study as shown in Tab. 4. We experiment with switching off $\mathcal{L}_{\text{depth}}$ and $\mathcal{L}_{\text{rigid}}$ – defined in Eq. 10 and Eq. 11 – together, as well as alternately. Omitting both losses leads to suboptimal accuracy. Including $\mathcal{L}_{\text{depth}}$ alone reduces the Abs Rel error from 0.121 to 0.103 and boosts the inlier percentage from 89.6% to 95.0%. The rigid consistency alone offers modest gains, but the best result is achieved by combining both losses, illustrating their mutual complementarity. Fig. 7 shows two intuitive comparisons between the reconstructions with and without our depth refinement. Directly fusing the monocular depth map with the estimated camera pose leads to inconsistent reconstructions with duplicated 3D structures, whereas our global refinement significantly improves 3D coherence.
Figure 7: Visualization of global refinement on the DAVIS [19]. We compare the 3D consistency of fused point clouds obtained from raw depth priors and our refined depths. Our refined depths yield a significantly more consistent 3D structure over time.
@@ 5 Conclusion
In this work, we addressed the challenge of reconstructing dynamic scenes from casual video sequences with a traditional SLAM method. We equipped a 3D point tracker with a motion decoupling mechanism to separate the camera-induced motion from object motion. Operating on the isolated static motion components, bundle adjustment leads to substantial improvements in the camera pose accuracy and reconstruction quality. Additionally, global refinement relies on the sparse depth estimates from BA to ensure scale consistency and alignment of depth maps across video frames. Overall, BA-Track leads to precise 3D trajectory estimation and dense reconstructions in highly dynamic environments. Our results show that deep priors can significantly enhance traditional optimization in challenging real-world settings.
@@ Appendix A Implementation Details
Model Architecture.   We develop our 3D tracker based on CoTracker [17], incorporating additional modifications. The feature extractor is a CNN-based architecture, consisting of a $7\times 7$ convolutional layer followed by several $3\times 3$ residual blocks, which generate multi-scale feature maps. These feature maps are aggregated into a single feature map with additional convolutional layers, producing a resolution of $\frac{1}{4}$ of the input image. We adopt the depth map encoding strategy from SpatialTracker [44], omitting the tri-plane correlation for improved efficiency. Instead, we introduce 2D depth correlation features during track updates to incorporate spatial information. The 3D tracker $\mathcal{T}$ is a transformer-based architecture that alternates between spatial and temporal attention blocks, consisting of 6 layers. For motion decoupling, we employ a smaller 3-layer transformer $\mathcal{T}_{\text{dyn}}$ designed specifically to predict the static component. Each layer in both transformers includes a temporal and a spatial attention block, with each block comprising an attention layer followed by an MLP.
Refinement Formulation.   The 3D tracker uses an iterative transformer-based refiner module $\mathcal{T}$ [17]. To provide the initial state as the input to $\mathcal{T}$, we copy the 2D location of the query point across all frames $s\in(1,...,S)$ and sample the point features as
$X^{(0)}(s)=X,\qquad f^{(0)}(s)=F_{t}[\mathbf{x}],$ (13)
@@
For dynamic tracker, $\mathcal{T}_{\text{dyn}}$ predicts the object-induced motion $X_{\text{dyn}}$ and dynamic point feature $f_{\text{dyn}}$ as
$(X_{\text{dyn}}^{(k+1)},f_{\text{dyn}}^{(k+1)})=\mathcal{T}_{\text{dyn}}(f_{\text{dyn}}^{(k)},\text{PE}(X_{\text{static}}^{(k)}),C_{f}^{(k)},C_{d}^{(k)}),$ (15)
@@
Local Context.   The local feature and depth correlation maps, $C_{f}^{(k)}$ and $C_{d}^{(k)}$, are functions of $X^{(k)}$, i.e.
$C_{f}^{(k)}(s)=C_{f}[X^{(k)}(s)],\quad C_{d}^{(k)}(s)=C_{d}[X^{(k)}(s)],$ (16)
and aim to enhance the spatial context of the point track. We compute $C_{f}$ and $C_{d}$ as follows. First, we generate multi-scale image feature maps $\tilde{\mathbf{F}}$ and depth maps $\tilde{\mathbf{D}}$ using bilinear interpolation from $\mathbf{F}$ and $\mathbf{D}$, respectively. Next, to compute the feature correlation for each point $C_{f}[X^{(k)}(s)]$, we take the inner product of the point feature $f^{(k)}(s)$ and the surrounding multi-scale image features $\tilde{\mathbf{F}}_{s}$ centered at $X^{(k)}(s)$ [17]. Similarly, we compute $C_{d}$ as the difference between the point’s depth component in $X^{(k)}(s)$, and that in the multi-scale local depth maps, $\tilde{\mathbf{D}}_{s}$, centered at $X^{(k)}(s)$.
@@
Training.   We train our model on the TAP-Vid-Kubric training set [10], which includes 11,000 sequences. Each sequence consists of 24 frames derived from the MOVi-F dataset. Following the data preprocessing steps from CoTracker [17], we additionally extract dynamic labels, 3D total trajectory and static trajectory ground truth. The static trajectory ground truth is generated by back-projecting queries from their 3D positions in the source frame into the target frames using ground truth camera poses. The original image resolution of each sequence is $512\times 512$, and we crop it to $384\times 512$ during training. For image augmentation, we apply random resizing, flipping, cropping, Gaussian blurring, and color jitter. For depth augmentation, we adopt the scale-shift augmentation and Gaussian blurring techniques described in  [14]. The augmented ground truth depth is consistently used during the training process.

### v2

@@ 3.3 Stage III: Global Refinement
Bundle adjustment recovers the camera poses and adjusts the depths of the query points in their source frames of reference. However, BA processes only a sparse set of query points, so the 3D positions of the other points in the depth maps remain unaffected. To ensure consistency of the depth maps with the sparse set of accurate 3D tracks estimated by BA, we need global refinement. Towards this goal, we introduce a function $\mathcal{H}_{\theta}:D\rightarrow\hat{D}$, where $\hat{D}$ is the refined depth map and $\theta$ is the parameter set. Global refinement jointly optimizes the function parameters $\theta$ and sparse 3D trajectories to achieve consistent geometry.
@@
Inspired by previous work [kopf2021robust], we define a 2D scale grid of resolution $H_{g}\times W_{g}$ for each frame, $\theta_{t}\in\mathbb{R}^{H_{g}\times W_{g}}$. The grid has a coarser resolution compared to the original image resolution $H\times W$. Its goal is to scale the depth estimates $D_{t}$: For each 2D query point $\mathbf{x}$ in frame $t$, we determine the associated depth scale using bilinear sampling on the grid $\theta_{t}$. The refined depth is then
@@
is the 3D position derived from back-projecting 2D points using the refined dense depth. The weight $W_{\text{static}}$ filters out the dynamic points as $W_{\text{static}}=(1-m^{i}_{a})\cdot(1-m^{i}_{b})$. Intuitively, $\mathcal{L}_{\text{rigid}}$ enforces constant distances between arbitrary static points $a$ and $b$ in the target frame $j$ and the source frame $i$. We optimize the variables using the Adam optimizer [kingma2014adam].
Figure 5: Qualitative camera pose estimation results on Sintel [Butler:2012:Sintel], Shibuya [qiu2022airdos], and Epic Fields [tschernezki2024epic]. Visualizations demonstrate that our method achieves more robust and accurate camera trajectories in challenging dynamic scenes.
@@ 4.1 Implementation Details
We train our 3D tracker on the TAP-Vid-Kubric training set [doersch2022tap], which consists of 11,000 sequences, each with 24 frames. The training process uses the AdamW optimizer [Loshchilov:2019:DWD] with a learning rate of $3\times 10^{-4}$ for a total of 100,000 steps, performed on 4 NVIDIA A100 GPUs. During training, we set the weights for visibility and dynamic labels to $w_{1}=5$ and $w_{2}=5$, employ $K=4$ iterative updates, a window size of $S=12$, and a total of $N=512$ query points. For BA, we use a window of 15 frames and apply 4 Gauss-Newton updates per window. We set $\alpha=0.05$ to balance reprojection and depth loss in Eq. 8. For BA weight filtering, we set $\delta_{v}=0.9$ and $\delta_{m}=0.9$. For further implementation details, please refer to our supplementary material.
@@ 4.2 Camera Pose Evaluation
Datasets.  We use three datasets for camera pose evaluation. MPI Sintel [Butler:2012:Sintel] provides sequences from 3D animated films featuring complex, fast object motion and exposure effects like motion blur. We use MPI Sintel to evaluate dynamic VO performance, with sequences spanning 20 to 50 frames. AirDOS Shibuya [qiu2022airdos] includes sequences of 100 frames tracking over 30 individuals in two scene types. “Road Crossing” features diverse object motion, while “Standing Humans” contains dominant camera motion. EPIC Fields [tschernezki2024epic] includes egocentric videos with dynamic interactions, such as moving hands with kitchen tools, and is therefore ideal for studying long dynamic scenarios. We randomly select 8 sequences with reliable ground-truth trajectories, sampling every third frame of the first 3000 frames, yielding 1000 frames per sequence.
Method MPI Sintel [Butler:2012:Sintel] AirDOS Shibuya [qiu2022airdos] Bonn [palazzolo2019iros]
@@
Robust-CVD [kopf2021robust] 0.703 47.8
CasualSAM [zhang2022structure] 0.387 54.7 0.261 79.9 0.169 73.7
MonST3R [ke2024repurposing] 0.335 58.5 (0.208) (71.2) 0.063 96.4
ZoeDepth [bhat2023zoedepth] 0.467 47.3 0.571 43.8 0.087 94.8
@@
test-time network finetuning; depth from image pairs; the numbers in $(\cdot)$ denote results on a sub-sequence due to memory constraints.
Table 2: Depth evaluation results on Sintel [Butler:2012:Sintel], Shibuya [qiu2022airdos], and Bonn [palazzolo2019iros] datasets. Our method achieves improved depth accuracy compared to the depth priors from ZoeDepth, demonstrating the effectiveness of our depth refinement.
Metrics.  We adopt three metrics. Absolute Translation Error (ATE) calculates the root mean square error between the estimated and ground-truth trajectories. Relative Translation Error (RTE) and Relative Rotation Error (RRE) measure translation (in meters) and rotation (in degrees) errors over a set distance, with both metrics averaged across all poses. We perform all evaluations after $\operatorname{Sim}(3)$ Umeyama alignment [umeyama1991least] with the ground-truth trajectory.
Results.  We compare BA-Track against learning-based SLAM and VO approaches, such as DROID-SLAM [teed2021droid], ParticleSfM [zhao2022particlesfm], DytanVO [shen2023dytanvo], and LEAP-VO [Chen:2024:LEAP]. Additionally, we compare BA-Track to test-time optimization techniques, such as Robust-CVD [kopf2021robust] and CasualSAM [zhang2022structure], which typically require several hours of processing time.
Table 1 shows a quantitative comparison of our method against state-of-the-art approaches across three benchmarks. Our approach consistently delivers superior performance in ATE and remains competitive in RTE and RRE. On the MPI Sintel dataset, our method significantly outperforms all baselines. As Fig. 5 illustrates, BA-Track produces notably more accurate trajectory estimates, particularly in dynamic scenes where other methods struggle. On the challenging Epic Fields dataset, characterized by rapid dynamic content and complex camera motion, BA-Track maintains compelling VO accuracy. Overall, BA-Track demonstrates remarkable accuracy across all benchmarks with high memory efficiency. By contrast, prior work like MonST3R can only process up to 90 frames at once on a 48GB GPU.
@@ 4.3 Depth Evaluation
Datasets.  MPI Sintel [Butler:2012:Sintel] and AirDOS Shibuya [qiu2022airdos] are selected as both provide dense ground-truth depth maps. We add a real-world dataset, Bonn RGB-D, which captures indoor activities such as object manipulation across 24 dynamic sequences. Following zhang2024monst3r, we select 5 dynamic sequences, each consisting of 110 frames.
Metrics.  We adopt the absolute relative error (Abs Rel) and the threshold accuracy (TA) $\delta<1.25$. Let $d_{i}$ and $d_{i}^{*}$ denote the predicted and the ground-truth depths for pixel $i$. Abs Rel is $\frac{1}{N}\sum_{i=1}^{N}\frac{|d_{i}^{*}-d_{i}|}{d_{i}^{*}}$ with $N$ the total number of pixels. TA is computed as the percentage of pixels satisfying $\max\left(\frac{d_{i}^{*}}{d_{i}},\frac{d_{i}}{d_{i}^{*}}\right)<1.25$. We apply shift-scale alignment by computing a single shift and scale factor over the entire video to align the predicted depths with the ground truth [zhang2024monst3r].
Figure 6: Motion decoupling on the DAVIS dataset [khoreva2019video]. From left to right, the three columns show the reference frame for queries, total observed motion, and the static component. Green indicates static trajectories, and red indicates dynamic ones. The static component retains points on dynamic objects in their original (static) positions within the corresponding reference frame.
Results.  We compare BA-Track to joint pose-depth optimization methods in Tab. 2. Using ZoeDepth [bhat2023zoedepth] as input, our method consistently enhances depth estimates and achieves competitive accuracy compared to other joint refinement methods. As we show later in Sec. 4.5, our global refinement plays a crucial role here, allowing BA-Track to handle challenging scenarios with dynamic content. On MPI Sintel, we achieve an Abs Rel of 0.408 and an accuracy of 54.1%, demonstrating comparable depth quality w.r.t. leading methods trained on much larger datasets, such as MonST3R [zhang2024monst3r]. In more complex scenarios, such as the outdoor urban scenes in AirDOS Shibuya, our method maintains robust results with an Abs Rel of 0.299 and 55.1% accuracy, outperforming competitors such as RobustCVD [kopf2021robust]. Our scale-map-based refinement offers efficient optimization with few parameters while remaining relatively simple. Investigating more advanced deformation models, such as neural networks, could be an interesting direction for future work.
@@ 4.4 Motion Decoupling
In our 3D tracker, motion decoupling isolates the static (camera-induced) component from the observed point motion, allowing us to treat dynamic points as if they were stationary. This approach reduces the impact of dynamic regions on camera pose estimation and dramatically enhances the robustness of BA. It also aids the subsequent depth refinement, providing more coherent reconstructions of both static and dynamic parts. Fig. 6 illustrates the observed and static point motion on the DAVIS [khoreva2019video] dataset. Red points indicate dynamic trajectories, while green points represent static ones, highlighting the 3D tracker’s ability to decouple the dynamic component from the observed motion.
To further demonstrate the benefits of motion decoupling, we analyze alternative motion representations for BA optimization in Tab. 3. By applying motion decoupling with the dynamic tracker $\mathcal{T}_{\text{dyn}}$, we achieve a dramatic performance boost: The ATE drops by nearly half, from 0.137 (a) to 0.065 (e). We observe a further moderate boost in VO accuracy by setting the weights of dynamic points in BA optimization to zero (f). Note that this surpasses the VO accuracy of the setting with a single tracker predicting the observed motion (b). As an alternative tracker architecture, we investigate an approach with a static tracker (c-d). In this case, we train the tracker to directly regress the static component. However, this approach leads to inferior VO accuracy compared to our decoupling approach (e-f). Presumably, a single network struggles to learn two aspects simultaneously: visual tracking and motion patterns. In contrast, our dual-network design addresses these tasks independently.
@@ 4.5 Dynamic Reconstruction
Table 3: Ablation study of dynamic handling methods on Sintel [Butler:2012:Sintel]. Combining motion decoupling with camera masking achieves the best results. * denotes a setting trained for ablation.
Method Depth Refinement Bonn crowd2 [palazzolo2019iros]
@@
Table 4: Ablation study on depth refinement on Bonn crowd2 sequence [palazzolo2019iros]. Applying both losses has a complementary effect.
We present qualitative results for camera tracking and 3D reconstruction on dynamic scenes in Fig. 2 using an alternative depth model from UniDepth [piccinelli2024unidepth], illustrating that our model is agnostic to the choice of depth priors. Using estimated camera poses, we back-project points into 3D via refined depth maps, yielding accurate camera motion estimation and consistent reconstructions. Our method remains robust in challenging scenarios with multiple moving people, rapid egocentric camera motion, and dynamic object interactions.
The superior reconstruction quality primarily stems from the proposed global refinement. To demonstrate this, we conduct an ablation study as shown in Tab. 4. We experiment with switching off $\mathcal{L}_{\text{depth}}$ and $\mathcal{L}_{\text{rigid}}$—defined in Eq. 10 and Eq. 11—together, as well as alternately. Omitting both losses leads to suboptimal accuracy. Including $\mathcal{L}_{\text{depth}}$ alone reduces the Abs Rel error from 0.121 to 0.103 and boosts the inlier percentage from 89.6% to 95.0%. The rigid consistency alone offers modest gains, but the best result is achieved by combining both losses, illustrating their mutual complementarity. Fig. 7 shows two intuitive comparisons between the reconstructions with and without our depth refinement. Directly fusing the monocular depth map with the estimated camera pose leads to inconsistent reconstructions with duplicated 3D structures, whereas our global refinement significantly improves 3D coherence.
Figure 7: Visualization of global refinement on the DAVIS [khoreva2019video]. We compare the 3D consistency of fused point clouds obtained from raw depth priors and our refined depths. Our refined depths yield a significantly more consistent 3D structure over time.
@@ 5 Conclusion
In this work, we addressed the challenge of reconstructing dynamic scenes from casual video sequences using a traditional SLAM method. We equipped a 3D point tracker with a motion decoupling mechanism to separate camera-induced motion from object motion. Operating on the isolated static motion components, bundle adjustment leads to substantial improvements in camera pose accuracy and reconstruction quality. Additionally, global refinement leverages the sparse depth estimates from BA to ensure scale consistency and alignment of depth maps across video frames. Overall, BA-Track enables precise 3D trajectory estimation and dense reconstructions in highly dynamic environments. More broadly, our work demonstrates that traditional optimization enhanced by deep priors can provide an accurate and robust solution to challenging real-world scenarios.
Acknowledgements.  This work was supported by the ERC Advanced Grant SIMULACRON, by the Federal Ministry for the Environment, Nature Conservation, Nuclear Safety and Consumer Protection (BMUV) through the AuSeSol-AI project (grant 67KI21007A), and by the TUM Georg Nemetschek Institute Artificial Intelligence for the Built World (GNI) through the AICC project.
[…]
@@ Appendix A Implementation Details
Model Architecture.  We develop our 3D tracker based on CoTracker [karaev2023cotracker], incorporating additional modifications. The feature extractor is a CNN-based architecture, consisting of a $7\times 7$ convolutional layer followed by several $3\times 3$ residual blocks, which generate multi-scale feature maps. These feature maps are aggregated into a single feature map with additional convolutional layers, producing a resolution of $\frac{1}{4}$ of the input image. We adopt the depth map encoding strategy from SpatialTracker [xiao2024spatialtracker], omitting the tri-plane correlation for improved efficiency.
The 3D tracker $\mathcal{T}$ is a transformer-based architecture that alternates between spatial and temporal attention blocks, consisting of 6 layers. For motion decoupling, we employ a smaller 3-layer transformer $\mathcal{T}_{\text{dyn}}$ designed specifically to predict the static component. Each layer in both transformers includes a temporal and a spatial attention block, with each block comprising an attention layer followed by an MLP.
Refinement Formulation.  The 3D tracker uses an iterative transformer-based refiner module $\mathcal{T}$ [karaev2023cotracker]. To provide the initial state as input to $\mathcal{T}$, we copy the 2D location of the query point across all frames $s\in(1,\dots,S)$ and sample the point features as
$X^{(0)}(s)=X,\hskip 28.80008ptf^{(0)}(s)=F_{t}[\mathbf{x}],$ (13)
@@
For the dynamic tracker, $\mathcal{T}_{\text{dyn}}$ predicts the object-induced motion $X_{\text{dyn}}$ and dynamic point feature $f_{\text{dyn}}$ as
$(X_{\text{dyn}}^{(k+1)},f_{\text{dyn}}^{(k+1)})=\mathcal{T}_{\text{dyn}}(f_{\text{dyn}}^{(k)},\text{PE}(X_{\text{static}}^{(k)}),C^{(k)}),$ (15)
@@
Local Context.  The local correlation map $C^{(k)}$ is a function of the point track $X^{(k)}$, i.e.
$C^{(k)}(s)=C[X^{(k)}(s)],$ (16)
and serves to enhance the spatial context of each tracked point. To compute $C$, we follow [xiao2024spatialtracker] and first apply Fourier embedding to the depth map $\mathbf{D}$ to obtain depth features $\mathbf{D}^{\text{Fourier}}$. These are concatenated with the image features $\mathbf{F}$ to produce fused features $\mathbf{F}^{\text{hyb}}$. We then apply bilinear interpolation to generate multi-scale hybrid feature maps $\tilde{\mathbf{F}}^{\text{hyb}}$. For each point $X^{(k)}(s)$, we extract a local region $\tilde{\mathbf{F}}^{\text{hyb}}_{s}$ centered at $X^{(k)}(s)$, and compute the correlation as the inner product between the point feature $f^{(k)}(s)$ and the surrounding hybrid features [karaev2023cotracker]. The resulting correlation map $C^{(k)}$ captures both appearance and geometric cues in a unified representation.
@@
Training.  We train our model on the TAP-Vid-Kubric training set [doersch2022tap], which includes 11,000 sequences. Each sequence consists of 24 frames derived from the MOVi-F dataset. Following the data preprocessing steps from CoTracker [karaev2023cotracker], we additionally extract dynamic labels, 3D total trajectory, and static trajectory ground truth. The static trajectory ground truth is generated by back-projecting queries from their 3D positions in the source frame into the target frames using ground-truth camera poses. The original image resolution of each sequence is $512\times 512$, and we crop it to $384\times 512$ during training. For image augmentation, we apply random resizing, flipping, cropping, Gaussian blurring, and color jitter. For depth augmentation, we adopt the scale-shift augmentation and Gaussian blurring techniques described in [gao2018ldso]. The augmented ground-truth depth is consistently used during the training process.